# A non-Hermitian loop for a quantum measurement

### v1

@@ Abstract
We establish a framework for a mechanism steering state vector collapse through time evolution as described by the Schrödinger’s equation. Under the premise of a non-Hermitian Hamiltonian dynamics, a main outcome is that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter. For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism effectively eliminating superpositions. Our results open a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
@@
Introduction and overview of key findings.– One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001). The crux of the measurement problem lies in explaining or making sense of the wave function collapse. The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see for example Cohen-Tannoudji et al. (2020); Shankar (1994)). This split description of the quantum dynamics is introduced to bridge the complex world of quantum superpositions with the familiar classical reality, where those superpositions have vanished. The incompatibility between the two descriptions for the dynamics entails a “cut” between the quantum and classical realms (as Werner Heisenberg called it in a posthumously published letter Bacciagaluppi and Crull (2011) in response to the EPR paper Einstein et al. (1935)). Indeed, the collapse is essentially a non-unitary, irreversible operation with loss of information and eventual energy and momentum exchange Pearle (2000), or as John Wheeler put it “an irreversible act of amplification such as the blackening of a grain of silver bromide emulsion” Wheeler (2016)  [^1] Wheeler mentions that the term irreversible amplification was used by Bohr, see chapter 1 of Bohr (2010), which was published in 1958..
@@
Despite advances in understanding the quantum-classical transition Zurek (2003); Schlosshauer (2007), the measurement problem remains a mystery for many of us Leggett (2006); Hance and Hossenfelder (2022)  [^2] There are also notable opinions against thinking of the measurement problem as such, see for example Mermin (2022). We also refer to Refs. Schlosshauer (2011); Schlosshauer et al. (2013) for a broader overview on the different positions.. Two proposals to address the conventional view of measurements are the Ghirardi-Rimini-Weber (GRW) Ghirardi et al. (1986, 1990) and the Continuous Spontaneous Localization (CSL) models Pearle (1989). Both introduce stochastic elements to the Schrödinger’s equation, suggesting that wave function collapse can occur spontaneously Bassi and Ghirardi (2003) (there are also other proposals using stochastic elements such as Fernández-Alcázar and Pastawski (2015)). Although GRW models do not rely on the concept of an apparatus to induce collapse, they still lack a microscopic explanation for how collapse occurs. Other proposals, inspired by GRW, are due to Diosi Diósi (1989); Donadi et al. (2021) and Penrose Penrose (1996) linking the wave function collapse to gravity. More recently, different authors (see  Martínez Romeral et al. (2024); Singh et al. (2023a, b) and also Van Wezel (2010); Mertens et al. (2023); Mukherjee and van Wezel (2024)) proposed using a non-Hermitian Hamiltonian to model a quantum measurement (earlier proposals also include dissipative modifications of the Schrödinger equation Gisin and Piron (1981)). These models are steps to make sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al. (2013) and the need for further experiments.
@@
In a parallel and unfolding story, over recent years, we have seen a rise in research fronts using photonics Rechtsman et al. (2013); Vicencio et al. (2015); Guzmán-Silva et al. (2014); Stützer et al. (2018), metamaterials Veenstra et al. (2024), acoustics Zhang et al. (2021), mechanical systems Lorenz et al. (2023) and circuits Lee et al. (2018); Imhof et al. (2018) as classical analogs for quantum systems. Notably, in photonic lattices, the paraxial Helmholtz equation for the electromagnetic waves maps to the Schrödinger equation Szameit and Nolte (2010) thus allowing to simulate quantum systems. Similarly, in electrical circuits, the correspondence between the circuit Laplacian and the tight-binding Hamiltonian have been extensively used to investigate topological states Lee et al. (2018); Imhof et al. (2018). Moreover, it has been demonstrated that these circuits can also mimic Schrödinger dynamics Ezawa (2019). Collectively, these examples illustrate the availability of a classical toolkit for quantum system simulation. However, a noteworthy limitation remains apparent: classical simulators have yet to fully address quantum measurements, which, because of the split description mentioned earlier, are distinct from quantum dynamics and thus remain a significant challenge.
In this Letter, we present a theoretical framework for non-Hermitian models of wave function collapse. Non-Hermitian Hamiltonians have been shown to exhibit many interesting features Moiseyev (2011); Longhi (2009); Rotter and Bird (2015). One of them is the appearance of non-Hermitian degeneracies, called exceptional points Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce. Exceptional points are at the heart of many counterintuitive Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012) and useful phenomena, such as the typical amplification of small parametric variations found near them which may lead to novel sensors  Wiersig (2014); Rechtsman (2017); Hodaei et al. (2017); Chen et al. (2017); Hokmabadi et al. (2019). Furthermore, non-Hermitian Hamiltonians offer a multitude of testing grounds in metamaterials Veenstra et al. (2024), photonics Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019); Vicencio et al. (2024), acoustics Zhang et al. (2021) and circuits Schindler et al. (2011); Hofmann et al. (2020), thereby offering a unique opportunity of simulating and discussing quantum collapse outside of the purely quantum domain.
@@
To set the stage, we first revisit the postulates of quantum mechanics, laying the groundwork for our discussion. Subsequently, we discuss a set of physically motivated assumptions and their consequences. The main outcome of this part is that, if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space. Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion: For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, the initial state undergoes a conversion to a preferred state, irrespective of initial conditions, which is determined by the sense of rotation in the space of parameters Uzdin et al. (2011); Berry and Uzdin (2011); Hassan et al. (2017a). Thus, chiral state conversion serves the purpose of breaking the superpositions and modeling quantum collapse under a single dynamic equation. However, there is a twist to consider: the sense of the loop is a property of the interaction itself, rather than the initial Hermitian Hamiltonian. In other words, it is an additional variable, and compatibility with standard quantum mechanics requires that its value must be set to satisfy Born’s rule. We close by discussing the implications of our results in the path to simulating a quantum measurement in classical systems.
@@
Split procedures for the quantum dynamics.– Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al. (2020); Shankar (1994). Two of these postulates specifically address the theory’s dynamics. One of them pertains to the standard, measurement-free behavior governed by the Schrödinger equation, highlighting how systems evolve when not being measured. The other postulate introduces the concept of measurement as a distinct phenomenon, where the act of measuring disrupts the usual dynamics, causing the system’s state vector to collapse into one of the observable’s eigenstates. This collapse results in the system adopting a state that corresponds to one of the observable’s eigenvalues, rendering the measurement as an exceptional process within quantum mechanics, described by a different antagonistic procedure.
@@
Figure 1: (a) Scheme showing a loop in a multidimensional parameter space defining the Hamiltonian. ${\cal H}_{0}$ is the initial Hamiltonian just before the interaction that leads to the collapse starts. (b) Time-evolution of the spin-up (blue) and spin-down (orange) relative probabilities for a system evolving under a Hamiltonian proportional to $\sigma_{x}$. A non-Hermitian term adding gains on the spin-up and loss on the spin-down is turned on with a gaussian function centered at time $t=5$ (standard deviation is $1$). This pulsed perturbation leads to a collapse after which the time-evolution restarts. (c) Scheme showing the Riemann surfaces of the real and imaginary part of the eigenvalues near an exceptional point. (d) Because of the peculiar shape of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
@@
Assumptions for a dynamical model of a quantum measurement.– Our discussion is based on the following set of assumptions for a family of models based on non-Hermitian Hamiltonians:
@@
    The measurement takes place over a finite time and the time evolution is dictated by the Schrödinger equation, with an eventual renormalization to keep the norm during the measurement process.
@@
    Over the timescale of the measurement, the Hamiltonian describes a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$, so that it remains unchanged at the end of the process (Fig. 1 (a)):
    $\begin{split}{\cal H}(\mathbf{X}(t=T))={\cal H}(\mathbf{X}(t=0))={\cal H}_{0}.\end{split}$ (1)
@@
    The interaction between the system and the measuring apparatus is modeled by introducing a non-Hermitian term to the Hamiltonian. For consistency with the previous points, this term needs to be time-dependent and describe a loop in parameter space. Details of the loop may also depend on the initial state.
Assumption 1 is the most natural one in order to recover unity in the description of the dynamics. Assumption 2 seems a reasonable assumption as we want for the system to remain the same when the apparatus (or the source of interaction) is removed. As a byproduct, the idea of making a closed loop in parameter space is naturally introduced.
Regarding assumption 3, the non-Hermitian term is a minimal model to account for the action of the measurement apparatus on the system without recurring to a larger Hilbert space, probably containing the degrees of freedom of the apparatus and/or the more complicated many-body interactions that escape the simple single-particle picture Foa Torres (2019). Another, more pragmatic motivation is that non-Hermitian Hamiltonians are readily available in quantum simulators using photonic lattices Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019) and electrical circuits Schindler et al. (2011); Hofmann et al. (2020).
All these assumptions can be summarized in two sentences: The measurement takes place over a finite time and is modeled as a non-Hermitian perturbation to an otherwise Hermitian Hamiltonian ${\mathcal{H}}_{0}$ which describes a loop in parameter space, thereby returning to the initial value after the interaction is complete. The parameter space is composed of those variables defining the Hamiltonian (including those related to the meter) and some of these parameters may have values set by the initial state.
The case of Ref. Martínez Romeral et al. (2024), where the collapse is obtained in a Stern-Gerlach setup (including spin and position variables), does not include an explicit time-dependence. However it can be thought as a particular equivalent case since its role is taken by the passage of the particle through the region where the magnetic field is acting.
Possible mechanisms.– Let us now turn to the discussion of the possible mechanisms. There are different ways where a non-Hermitian term could lead to a collapse of a superposition state on one of its components. A very simple way would be just adding gains and losses as to amplify one of the possible states and reduce the others. A typical case is shown in Fig.1 (b), where a superposition state is collapsed onto one of its components. In this case, we take for illustration purposes:

### v2

@@ Abstract
Here we present a non-Hermitian framework for modeling state-vector collapse under unified dynamics described by Schrödinger’s equation. Under the premise of non-Hermitian Hamiltonian dynamics, we argue that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter. For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism for effectively eliminating superpositions. This perspective opens a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
I Introduction
@@
One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001). The crux of the measurement problem lies in explaining or making sense of the wave function collapse. The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see, for example, Cohen-Tannoudji et al. (2020); Shankar (1994)). This split description of the quantum dynamics is introduced to bridge the complex world of quantum superpositions with the familiar classical reality, where those superpositions have vanished. The incompatibility between the two descriptions for the dynamics entails a “cut” between the quantum and classical realms (as Werner Heisenberg called it in a posthumously published letter Bacciagaluppi and Crull (2011) in response to the EPR paper Einstein et al. (1935)). Indeed, the collapse is essentially a non-unitary, irreversible operation with loss of information and eventual energy and momentum exchange Pearle (2000), or as John Wheeler put it “an irreversible act of amplification such as the blackening of a grain of silver bromide emulsion” Wheeler (2016)  [^1] Wheeler mentions that the term irreversible amplification was used by Bohr, see chapter 1 of Bohr (2010), which was published in 1958..
@@
Despite advances in understanding the quantum-classical transition Zurek (2003); Schlosshauer (2007), the measurement problem remains a mystery for many of us Leggett (2006); Hance and Hossenfelder (2022)  [^2] There are also notable opinions against thinking of the measurement problem as such, see for example Mermin (2022). We also refer to Schlosshauer (2011) for a broader overview of the different positions. Two proposals to address the conventional view of measurements are the Ghirardi-Rimini-Weber (GRW) Ghirardi et al. (1986, 1990) and the Continuous Spontaneous Localization (CSL) models Pearle (1989). Both introduce stochastic elements to the Schrödinger’s equation, suggesting that wave function collapse can occur spontaneously Bassi and Ghirardi (2003) (there are also other proposals using stochastic elements such as Fernández-Alcázar and Pastawski (2015)). Although GRW models do not rely on the concept of an apparatus to induce collapse, they still lack a microscopic explanation for how collapse occurs. Other proposals, inspired by GRW, are due to Diósi Diósi (1989); Donadi et al. (2021) and Penrose Penrose (1996) linking the wave function collapse to gravity. More recently, different authors (see  Martínez Romeral et al. (2024); Singh et al. (2023a, b) and also Van Wezel (2010); Mertens et al. (2023); Mukherjee and van Wezel (2024)) proposed using a non-Hermitian Hamiltonian to model a quantum measurement (earlier proposals also include dissipative modifications of the Schrödinger equation Gisin and Piron (1981)). These models are steps in making sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al. (2013) and the need for further experiments.
@@
In a parallel and unfolding story, over recent years, we have seen a rise in research fronts using photonics Rechtsman et al. (2013); Vicencio et al. (2015), metamaterials Veenstra et al. (2024), acoustics Zhang et al. (2021), mechanical systems Lorenz et al. (2023) and circuits Lee et al. (2018) as classical analogs for quantum systems. Notably, in photonic lattices, the paraxial Helmholtz equation for the electromagnetic waves maps to the Schrödinger equation Szameit and Nolte (2010) thus allowing to simulate quantum systems. Similarly, in electrical circuits, the correspondence between the circuit Laplacian and the tight-binding Hamiltonian have been extensively used to investigate topological states Lee et al. (2018). Moreover, it has been demonstrated that these circuits can also mimic Schrödinger dynamics Ezawa (2019). Collectively, these examples illustrate the availability of a classical toolkit for simulation of quantum systems. However, a critical limitation remains: while a few noteworthy exceptions offer solutions for specific cases Longhi (2006); Liu et al. (2023), classical simulators generally lack a broadly applicable counterpart to quantum measurements, thereby preventing the exploration of a wealth of phenomena such as measurement-enriched phases.
The aforementioned limitation of classical simulators in handling quantum measurement shows the need for a deeper understanding of the measurement process itself, which might then inspire new simulation approaches. Focusing on recent proposals modeling collapse using non-Hermitian Hamiltonians, challenges remain. Although models like our prior work Martínez Romeral et al. (2024) show how non-Hermitian dynamics can amplify a chosen state, the need to construct the interaction specifically to favor a predetermined outcome perhaps offers less insight into the measurement process itself. Moreover, such approaches do not necessarily provide a dynamical explanation for why the possible outcomes of an energy measurement correspond specifically to the eigenstates of the system’s original Hamiltonian ($\mathcal{H}_{0}$). One might desire a mechanism where both the preference for eigenstates and the selection among them emerge dynamically from the system’s response to control or interaction. This motivates our exploration of an alternative framework centered on dynamically induced collapse via parameter loops around exceptional points, aiming to demonstrate how such dynamics can naturally establish the initial eigenstates as the unique stable outcomes and provide a topological mechanism for selecting one.
The alternative framework we propose leverages key features of non-Hermitian Hamiltonians Moiseyev (2011); Bergholtz et al. (2021); Lin et al. (2023), systems whose unique properties offer a potential pathway to address the limitations discussed above. Non-Hermitian Hamiltonians exhibit phenomena absent in their Hermitian counterparts, most notably non-Hermitian degeneracies known as exceptional points (EPs) Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce. These EPs are central to many counterintuitive phenomena Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012) and useful effects, including sensitivity enhancements relevant for novel sensors Wiersig (2014); Rechtsman (2017); Hodaei et al. (2017); Chen et al. (2017); Hokmabadi et al. (2019). Crucially for our goal of dynamically explaining state selection, the behavior near EPs, particularly the state evolution during parameter loops encircling them Hassan et al. (2017a), forms the core of our proposed collapse mechanism. Furthermore, the physics of non-Hermitian systems and EPs can be readily explored in various experimental platforms, including metamaterials Veenstra et al. (2024), photonics Szameit et al. (2011); Pan et al. (2018); Miri and Alù (2019); Vicencio et al. (2025), acoustics Zhang et al. (2021), and circuits Schindler et al. (2011); Hofmann et al. (2020). This experimental accessibility offers a unique opportunity not only to test the fundamental dynamics we propose but also potentially to simulate aspects of quantum collapse, as modeled here, within controllable classical systems.
@@
To set the stage, we first revisit the postulates of quantum mechanics, laying the groundwork for our discussion. Subsequently, we discuss a set of physically motivated assumptions and their consequences. The main outcome of this part is that if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space (returning to its initial Hermitian form). Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion Hassan et al. (2017a): For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, and under certain conditions for the speed with which the loop is completed, the initial state undergoes a conversion to a preferred state, which is determined by the sense of rotation in the space of parameters Hassan et al. (2017a). Thus, chiral state conversion breaks the superpositions and models quantum collapse under a single dynamic equation. However, there is a twist to consider: the sense of the loop is a property of the interaction itself rather than the initial Hermitian Hamiltonian. In other words, it is an additional variable, and compatibility with standard quantum mechanics requires that its value must be set to satisfy Born’s rule. We close by discussing the implications of our results on the path to simulating quantum measurement in classical systems.
II Split procedures for the quantum dynamics
@@
Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al. (2020); Shankar (1994). Two of these postulates specifically address the theory’s dynamics. One of them pertains to the standard, measurement-free behavior governed by the Schrödinger equation, highlighting how systems evolve when not being measured. The other postulate introduces the concept of measurement as a distinct phenomenon, where the act of measuring disrupts the usual dynamics, causing the system’s state vector to collapse into one of the observable’s eigenstates. This collapse results in the system adopting a state that corresponds to one of the observable’s eigenvalues, rendering the measurement as an exceptional process within quantum mechanics, described by a different antagonistic procedure.
The measurement postulate in quantum mechanics, in its conventional form, can be broken down into three key questions: 1) What are the possible outcomes of a measurement? 2) What happens to the system’s state after a measurement (state-vector collapse)? and 3) What are the probabilities of obtaining each outcome (Born’s rule)? Here, we primarily address the first and second question, offering a model for the dynamics of the collapse itself.
@@
Figure 1: (a) Scheme showing a loop in a multidimensional parameter space defining the Hamiltonian. ${\cal H}_{0}$ is the initial Hamiltonian just before the interaction that leads to the collapse starts. (b) Time-evolution of the spin-up (blue, continous) and spin-down (orange, dashed) relative probabilities for a system evolving under the Hamiltonian of Eq.  2 with $\gamma_{0}=1$ and $\Gamma=50$. The state at $t=0$ corresponds to a particle with spin down. A non-Hermitian term adding gains on the spin-up and loss on the spin-down is turned on with a gaussian function centered at time $t=5$ (standard deviation is $1$). This pulsed perturbation leads to a collapse after which the time-evolution restarts. (c) Scheme showing the Riemann surfaces of the real and imaginary part of the eigenvalues near an exceptional point. (d) Because of the peculiar structure of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
@@
Dynamical Model of a Quantum Measurement.– Our model for a dynamical description of quantum measurement using non-Hermitian Hamiltonians rests on three key assumptions:
@@
    Finite-Time Evolution: The measurement occurs over a finite time interval. The system’s state vector, denoted as $|\Psi(t)\rangle$, evolves according to the standard linear time-dependent Schrödinger equation:
    $i\hbar\frac{\partial|\Psi(t)\rangle}{\partial t}=\mathcal{H}(t)|\Psi(t)\rangle,$
    where $\mathcal{H}(t)$ is the (generally non-Hermitian) Hamiltonian encoding the system and its interaction with the measurement apparatus. The resulting state vector $|\Psi(t)\rangle$ is not necessarily normalized to unity; its components (i.e., projections onto a basis, such as $\langle\text{outcome}|\Psi(t)\rangle$) are interpreted as relative probability amplitudes. To obtain physical probabilities for specific outcomes at any given time $t$, the squared magnitudes of these relative probability amplitudes (yielding relative probabilities) are subsequently normalized by the total instantaneous norm $\langle\Psi(t)|\Psi(t)\rangle$ (i.e., $P_{\text{outcome}}(t)=|\langle\text{outcome}|\Psi(t)\rangle|^{2}/\langle\Psi(t)|\Psi(t)\rangle$). We note that alternative theoretical approaches exist where norm preservation is enforced continuously via non-linear modifications to the Schrödinger equation (see, e.g., Ref. Brody and Graefe (2012)).
@@
    Closed Loop in Parameter Space: The Hamiltonian $\mathcal{H}(\mathbf{X}(t))$ traces a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$ during the measurement, ensuring that it returns to its initial form $\mathcal{H}_{0}$ by the end of the process (Fig. 1(a)). This can be represented as:
    $\mathcal{H}(\mathbf{X}(t=T))=\mathcal{H}(\mathbf{X}(t=0))=\mathcal{H}_{0}$ (1)
@@
    Non-Hermitian term: The interaction with the measurement apparatus is modeled by a time-dependent non-Hermitian term in the Hamiltonian. This term also describes a loop in parameter space, potentially dependent on the initial state of the system.
These assumptions ensure that the system returns to its original form after the measurement apparatus is removed. The non-Hermitian term serves as a minimal model for the apparatus’s influence, avoiding the need to explicitly include the apparatus’s degrees of freedom in the Hilbert space, which would lead to a more complex many-body problem, and is readily implemented in quantum simulators such as photonic lattices Szameit et al. (2011); Pan et al. (2018); Miri and Alù (2019) and electrical circuits Schindler et al. (2011); Hofmann et al. (2020). The case of Ref. Martínez Romeral et al. (2024), where collapse is obtained in a Stern-Gerlach setup including both spin and position variables does not include an implicit change in time. However, it can be viewed as a particular case where the particle’s passage through the magnetic field region effectively introduces the time dependence when viewed on the spin variable.
In essence, our model describes the measurement-induced collapse as resulting from a continuous (and not instantaneous) process, driven by a non-Hermitian perturbation on the initial Hermitian Hamiltonian, which could be mapped to a close loop dynamics in parameter space, dictated by the variables of the Hamiltonian, some of which may be influenced by the initial state.
III Possible mechanisms
Let us now turn to the discussion of the possible mechanisms. There are different ways where a non-Hermitian term could lead to a collapse of a superposition state on one of its components. A very simple way would be just to add gains and losses to amplify a target state and reduce the others. A typical case is shown in Fig.1 (b), where a superposition state is collapsed onto one of its components. In this case, we take for illustration purposes: